If the probability that a certain tennis player will serve an ace is 1/4, what is the probability that he will serve exactly two aces out of four serves? (Assume that the four serves are independent. Round your answer to four decimal places.)

Answer by stanbon(75887)

If the probability that a certain tennis player will serve an ace is 1/4, what is the probability that he will serve exactly two aces out of four serves?
-----
Binomial Problem with n = 4 and p(ace) = 1/4
----
P(x = 2) = 4C2*(1/4)^2*(3/4)^2 = binompdf(4,1/4,2) = 0.2109
=======================
Cheers,
Stan H.
=======================